Tiggie’s Dog Toys, Inc. reported a debt-to-equity ratio of 1.75 times at the end of 2018. If the firm’s total assets at year-end were $25 million.

How much of their assets are financed with debt and how much with equity? (Do not round intermediate calculations. Enter your answer in millions of dollars rounded to 3 decimal places.)
Answer #1

Debt/Equity=1.75
=>(Debt+Equity)/Equity=(1+1.75)/1=2.75/1
=>Equity/(Debt+Equity)=1/2.75
Weight of equity=1/2.75

Now, 1-Equity/(Debt+Equity)=1-1/2.75
=>Debt/(Debt+Equity)=1.75/2.75
Weight of debt=1.75/2.75

Total assets=$25 million

Assets are financed with equity=Weight of equity*Total assets=1/2.75*25=9.090909091 or $9.091 million (Rounded to 3 decimal places)
Assets are financed with debt=Weight of debt*Total assets=1.75/2.75*25=$15.909 million